Milatin Group reported the following revenues and net income amounts (in millions) Revenue Net income 2017 2016 2015 2014 13,200 $11,520 S 1,040 12,000 $567 $ 513 S 486 S 540 Show Milatin Groups trend percentages for revenues and net income. Use 2014 as the base year and round to the nearest percent. Which measure increased faster during 2015-2017? a. b. a. Show Milatin Groups trend percentages for revenues and net income. Use 2014 as the base year and round to the nearest percent. (Round your answers to the nearest whole percent.) Trend percentages Revenue Net income b. Which measure increased faster during 2015-2017? During 2015 to 2017 2017 2016 2015 2014 100 % 100 % ▼ increased faster than revenue net income
Answer #1

a) Trend percentage

Trend percentage 2017 2016 2015 2014
Revenue 13200/12000 = 110% 11520/12000 = 96% 11040/12000 = 92% 100%
Net income 567/540 = 105% 513/540 = 95% 486/540 = 90% 100%

During 2016 to 2017 Revenue increased faster than Net income